Traveling with the current, a boat covers in 4 hours a distance 2.4 times greater than traveling against the current. What is the speed of the boat in still water if the speed of the current is 1.5 mph?

Answer by ankor@dixie-net.com(22740)

Traveling with the current, a boat covers in 4 hours a distance 2.4 times greater than traveling against the current.
What is the speed of the boat in still water if the speed of the current is 1.5 mph?
:
let s = boat speed in still water
then
(s+1.5) = speed with the current
and
(s-1.5) = speed against
:
Write a distance equation, dist = time * speed
:
with dist = 2.4*against dist
4(s+1.5) = 2.4[4(s-1.5)]
4s + 6 = 2.4(4s - 6)
4s + 6 = 9.6s - 14.4
6 + 14.4 = 9.6s - 4s
20.4 = 5.6s
s = 20.4/5.6
s = 3.643 mph in still water
:
;
See if that checks out
4(3.643+1.5) = 2.4[4(3.643-1.5)]
4(5.143) = 2.4*4(2.143)
20.57 = 2.4(8.57)
20.57 = 20.57